OpenStudy (anonymous):

identify the missing factor (3^2a^3)(?) = -18a^6

OpenStudy (gw2011):

Let x = Missing factor (3)^2(a^3)(x) = -18a^6 (9a^3)(x) = -18a^6 x = (-18a^6)/9a^3 x = -2a^3 This is the missing factor